OpenStudy (anonymous):

graph the linear system and estimate the solution. then check the solution algebraically. x+2y=2 x-4y=14

OpenStudy (anonymous):

how do we graph it? do we make it to y=-1/2x and y=1/4x-3.5 first?

OpenStudy (precal):